OpenStudy (anonymous):

what percent of 13.76 is 86?

OpenStudy (anonymous):

13.76*(x/100)=86 x=8600/13.76=625

OpenStudy (anonymous):

thank you so muchhh!!

OpenStudy (anonymous):

if i deserve medal then give me

OpenStudy (anonymous):

u got it lol

OpenStudy (anonymous):

thanx..........any more q

OpenStudy (anonymous):

yea h/o

OpenStudy (anonymous):

-3( x+1) =x-5+2(5x+5)